A customer desires to mix her own concrete. The clerk knows that the ratio of cement to sand to gravel is 1:3:4. If the customer needs 1,000 pounds of dry mix, how many pounds of each ingredient should she use?
Please use cross products

Answer by checkley77(12844)

The direct way to solve this problem is to add the three ratios:
1+3+4=8
Now find how many sets of 8 there are in the total of 1,000.
1,000/8=125 for each 1 of the elements in the ratio.
1*125=125 pds. of cement.
3*125=375 pds. of sand.
4*125=500 pds. of gravel.
Proof:
125+375+500=1,000
1,000=1,000